# Statistical Inference for Extended Target Detection in mmWave Automotive Radar

### v2

@@ Abstract
Millimeter-wave (mmWave) automotive radar systems offer high range resolution due to their wide bandwidth, enabling the detection of multiple spatially distributed scatterers from a single extended target, such as a vehicle. Traditional CFAR-based detection methods often treat these scatterers as independent point targets, thereby neglecting the inherent spatial structure of extended objects. To address this limitation, we propose a novel Range-Doppler (RD) segment-based statistical inference framework that captures the characteristic scattering profile of extended automotive targets. The framework employs Maximum Likelihood Estimation (MLE) for statistical parameter extraction and utilizes Gibbs sampling within a Markov Chain Monte Carlo (MCMC) scheme to model the posterior distribution of the segment features. A skewness-based test statistic, derived from the estimated distribution, is introduced for binary hypothesis testing to distinguish extended targets. Furthermore, we develop a detection pipeline incorporating Intersection over Union (IoU) metrics and peak-centric segment alignment, optimized for single-dwell radar operations. Comprehensive evaluations on both simulated and real-world datasets demonstrate the proposed method’s effectiveness, achieving enhanced detection accuracy and robustness in automotive radar scenarios.

### v3

# Gamma-Based Statistical Modeling for Extended Target Detection in mmWave Automotive Radar

@@ Abstract
Millimeter-wave (mmWave) radar systems, owing to their large bandwidth, provide fine range resolution that enables the observation of multiple scatterers originating from a single automotive target—commonly referred to as an extended target. Conventional CFAR-based detection algorithms typically treat these scatterers as independent detections, thereby discarding the spatial scattering structure intrinsic to the target. To preserve this scattering spread, this paper proposes a Range-Doppler (RD) segment framework designed to encapsulate the typical scattering profile of an automobile. The statistical characterization of the segment is performed using Maximum Likelihood Estimation (MLE) and posterior density modeling based on the Gamma distribution, facilitated through Gibbs Markov Chain Monte Carlo (MCMC) sampling.A skewness-based test statistic, derived from the estimated statistical model, is introduced for binary hypothesis classification of extended targets. Additionally, the paper presents a detection pipeline that incorporates Intersection over Union (IoU) and segment centering based on peak response, optimized to work within a single dwell. Extensive evaluations using both simulated and real-world datasets demonstrate the effectiveness of the proposed approach, underscoring its suitability for automotive radar applications through improved detection accuracy.